OpenStudy (anonymous):

Bank of America customers select their own three-digit personal identification number (PIN) for use at ATMs. (b) What is the probability Mr. Jones and Mrs. Smith select the same PIN? (c) Which concept of probability did you use to answer (b)? Justify your answer

OpenStudy (alexwee123):

there are 1000 possible pins so what do you think the answer is?

OpenStudy (anonymous):

dun knw....

OpenStudy (shane_b):

Since you only have 3 digits available for PINs the range would be 000-999, which is 1000 possible pins as @alexwee123 said. For two people to select the same pin, it would be:\[\frac{1}{1000}*\frac{1}{1000}=\frac{1}{1000000}\]So 1 in a million...

OpenStudy (anonymous):

thx....how about (c)

OpenStudy (shane_b):

I have no idea of what the concept is called...I just remember the math :)

OpenStudy (anonymous):

ok....thx...